Bob needs to buy 5 pounds of candy. one kind costs 1.75 per pound and the second kind costs 2.35 per pound. How much of each will he need if he spends 10.00?
Answer by ankor@dixie-net.com(22740)

Bob needs to buy 5 pounds of candy. one kind costs 1.75 per pound and the second kind costs 2.35 per pound. How much of each will he need if he spends 10.00?
:
A mixture problem;
:
We can say that the resulting mixture is worth 10/5 = 2.00/lb
:
let x = amt of 1.75 /lb candy;
let (5-x) = 2.35 /lb candy
:
1.75x + 2.35(5-x) = 2(5)
:
1.75x + 11.75 - 2.35x = 10
1.75x - 2.35x = 10 - 11.75
-.6x = -1.75
x = -1.75/-.6
x = 2.92 lb of the $1.75 candy
:
5 - 2.92 = 2.08 lb of the $2.35 candy
:
:
Check:
1.75(2.92) + 2.35(2.08) =
5.11 + 4.89 = $10